LgswUnitDx work

8:30: WG and MR go to check the cal unit illumination patter. It seems ok, maybe an alignment problem?

10:20: BcuNumber3 fails to connect. We must power cicle it 3 times to get it working...

12:20: Still problems of shutdown with the BCU. We connect the power line to the oscilloscope. Basically when the BCU disconnects the voltage goes to 0... This happens between 2 and 50min from configuration, even if the BCU is left idle (no sloping, no jittering).

We have also lot of problems with the AdSecDx. It is very hard to get it back after a failure... Aslo when we enable the slopes from our BCU the shell rips...

13:30: We setup BcuNumber2 for replacing the actual DX.

Finally we succeed in closing a loop with May 15 configuration and calibrations...

WFS PM X PM Y
Blue 420 530
Yellow 280 390
Red 405 360

TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20161016_172500 0, 0, 0 F NA NA NA -1 F F
20161016_215239 0.2,0.1,0.05 T 224.265 70.5766 20150704_021200 -1 F F
16:30: We swap the BCUs, but HVC board 0010 has a problem the channel 0 seems broken
Topic revision: r2 - 20 Oct 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback